Circular No. 7392

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000bg IN NGC 6240 Y. Sato and W. D. Li, University of California at Berkeley, on behalf of

the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Apr. 1.5 and 2.5 UT, both with mag ∼ 17.4. SN 2000bg is located at α = 16^{h}52^{m}58^{s}.13, δ = +2^{o}23′50′′.5 (equinox 2000.0), which is 11′′.4 west and 12′′.0 south of the nucleus of NGC 6240. An unfiltered image taken on Mar. 23.5 showed nothing at the position of the new object (limiting mag ∼ 19.0).

XTE J1118+480 M. Garcia, W. Brown, M. Pahre, and J. McClintock, Harvard-Smith-

sonian Center for Astrophysics; P. Callanan, University College, Cork; and P. Garnavich, University of Notre Dame, report: “Optical spectroscopy of XTE J1118+480 on Mar. 31 UT reveals a spectrum typical of an x-ray nova in outburst (e.g., Shrader et al. 1994, Ap.J. 434, 698). A series of spectra obtained with the F.L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph) show the following emissions, equivalent widths, and FWHM (accurate to 10 percent): Hα, 0.19 nm, 2200 km/s; Hβ, 0.01, –; He II 468.6-nm, 0.12, 2690. The Hα FWHM that we find in this x-ray nova is approximately equal to the largest seen in quiescent blackhole x-ray novae. The Hα line shows two blueward absorption features, and the Hβ line shows redward and blueward absorption features. The Na D 589.0-nm line has an equivalent width of 0.004 nm, indicating very low interstellar absorption. The column density out of the plane in this direction (Dickey and Lockman 1990, Ann. Rev. Astron. Ap. 28, 215) corresponds to E(B − V) ∼ 0.024. Differential V -band photometry from the FLWO 1.2-m telescope, relative to the USNO-A2.0 star at α = 11^{h}18^{m}40^{s}.65, δ = +48^{o}03′25′′.8 (equinox 2000.0; V = 10.91) yields magnitudes of 12.80(0.1) and 13.05(0.1) on Mar. 31.151 and 31.396, respectively. The surprisingly low x-ray-to-optical flux ratio (i.e., V ∼ 13, F_{x} ∼ 40 mCrab; cf. IAUC 7389) may be due to a nearly edge-on viewing angle.”

G. Masi, Ceccano, Italy, provides the following position for the optical counterpart to XTE J1118+480 (cf. IAUC 7390) from ten co-added 25s unfiltered CCD images: α = 11^{h}18^{m}10^{s}.85, δ = +48^{o}02′12′′.9 (equinox 2000.0; twelve USNO-A2.0 reference stars, mean residual 0′′.2); the star’s magnitude was 13.0 on Apr. 1.825 UT (limiting mag ∼ 18).

2000 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green